a rectangular swimming pool has a diagonal length of square root 68 meters and a length that is 6 more meters than width. find the dimensions and area of an entire pool
Answer by fractalier(6550)

Let's call its width, x. That means the length is x+6.
Now we apply the Pythagorean Theorem:
a^2 + b^2 = c^2
x^2 + (x+6)^2 = (sqrt(68))^2
x^2 + x^2 + 12x + 36 = 68
2x^2 + 12x - 32 = 0
x^2 + 6x - 16 = 0
Now factor and solve
(x + 8)(x - 2) = 0
x = -8 (not possible)
or
x = 2 meters (width) so that the length is
x+6 = 8 meters